Shift report 20000822 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Aug 22 2000 - 08:00:42 EDT

  • Next message: Konstantin Olchanski: "operator account emergency cleanup"

    ------------------------------------------------------------
    Date: Tue Aug 22 08:00:42 2000, Shift: 00:00-08:00
    Supervisor: YB, with: B1
    Log book#4 pages 138 - 140
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was on since midnight. At 5:30am it was dumped by 
       mutual request of several experiments. Beam is back since
       6:30am.
    2. Runs 2379 -- 2384 is still running. Total of about 450K 
       events (trig 1).
    3. T1 was tripping every .5 hour or so ==> turned it off.
       Had trouble to turn on the magnets, so now there's a 
       short manual in the "BRAHMS Shift Leaders and Operations"
       book on the magnet page.
       "operator" account ran out of space. Moved all old 
       run0014*.root files to /pii3/scratch/ntuple/ 
       "operator" account has to be cleaned and/or new place for 
       "*.root" files has to be found. BrNtuple can't handle 
       runs with a lot of sequences (more than 12-15?) Created 
       several ntuples for the same run: "run237900.root" has
       run002379seq000 -- 009.dat files in it, and so on. 
       Couldn't do the rest of the runs (no space!).
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Aug 22 2000 - 08:05:03 EDT